OpenStudy (anonymous):

The square root of 5 (the square root of 5 +3) Simplify please? (:

OpenStudy (anonymous):

i dont understand the question ...u want root 5 or root 8

OpenStudy (anonymous):

\[\sqrt{5} (\sqrt{5}+\sqrt{3})\]

OpenStudy (anonymous):

oh okk...multiply it inside.√a(√a+√b) = √a*√a+√a*√b = a + √ab

OpenStudy (anonymous):

That clear?

OpenStudy (anonymous):

What?? What's up the the a's and b's?! Where'd you get that??

OpenStudy (anonymous):

lol okk ...√5(√5+√3) = √5*√5+√5*√3 = 5 + √15

OpenStudy (anonymous):

clear now?

OpenStudy (anonymous):

So.. What's the answer? I'm like SO CONFUSED.

OpenStudy (anonymous):

5 + √15

OpenStudy (anonymous):

Oh okay. Thank you(:

OpenStudy (anonymous):

You're welcome! :)